Compute the takt time for a system where the total time per shift is 480 minutes, there are two shifts, and workers on each shift are given two 15-minute breaks and 45 minutes for lunch. Daily demand is 650 units.

What cycle time would match capacity and demand if demand is 120 units a day, there are two shifts of 480 minutes each, and workers are given two half-hour breaks during each shift, one of which is for lunch or dinner?

Compute the takt time for a service system that has a total work time of 440 minutes per day, two 10-minute breaks, and an hour for lunch. The service system must process 100 jobs per day.
Answer #1

1. available time per shift = 480 - (15x2+45) = 405

Total time available per day = 405x2 =810

takt time = total time available / number of units to be produced = 810/650 = 1.246 min.

2. Cycle time needed to meet the demand = 2x[480-2x30] / 120 = 7 min per unit

3. takt time = 440-(10x2+60) /100 = 3.6 min.